### v4

# Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed

Abstract.
In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live stream policy for accurate live stream allocation. The inappropriate live stream allocation policy can significantly affect the duration of the usage app and user retention, which ignores the long-term negative impact of live stream allocation. Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement. However, traditional RL algorithms often face divergence and instability problems, which restricts the application and deployment in the large-scale industrial recommendation systems, especially in the aforementioned challenging scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods under the platform-level constraints but also exhibits enhanced stability in online recommendation scenarios.
Reinforcement Learning, Recommendation System, Variance Reduction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
Figure 1. Structure of a short video & live stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live stream RS.
1. Introduction
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
Currently, deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
Furthermore, like advertisement, live stream is not welcomed by every user. Allocating too many live streams for a user will significantly interrupt the user’s short video interest and lead to a decrease in the user’s app usage duration, which will eventually affect the user engagement and user retention. Therefore, we cannot merely focus on the user feedback for a single request, but rather should optimize the long-term user experience.
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
The main contributions of this paper are as follows.
  • We introduce SL-MGAC, a novel RL model that outperforms existing methods in maximizing the user engagement of live stream while satisfying the platform-level constraints on app usage duration and user retention.
  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q value normalization.
  • We integrate multi-task supervised learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
  • We successfully deploy the SL-MGAC model in a challenging short video and live stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
2. Problem Formulation
As shown in Fig. 1, in the short video & live stream mixed recommendation system, the live stream recommendation system is treated as an agent that interacts with various users and receives user feedback over time. However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention. Therefore, the optimal live stream allocation control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
Formally, we define live stream allocation CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count). We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request. We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream and $a=0$ means not injecting it.
  • Transition Probability $\mathcal{P}$: The transition probability is denoted as $p(s_{t+1}|s_{t},a_{t})$, determined by the environment.
  • Reward Function $\mathcal{R}$: The reward function $\mathcal{R}$ is a mapping from state $s_{t}$ and action $a_{t}$ at timestamp $t$, which can be formulated as $r(s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the calculation of the cumulative reward. Typically, we set $\gamma<1$ during model training.
  • Constraint $\mathcal{C}$: The platform-level constraint is an instantaneous constraint, that is, the live stream watch time should be longer than the average video watch time per request as much as possible. This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter. By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
Figure 2. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
Overall, the concrete reward function $r(s_{t},a_{t})$ and constraint function $c(s_{t},a_{t})$ is shown below:
(1) $r(s_{t},a_{t})$ $=y_{l}$
$c(s_{t},a_{t})$ $=\frac{1}{B}y_{v}-y_{l}$
where $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
The optimization objective of aforementioned optimal live stream allocation control problem is shown as follows:
(2) $\max_{\pi}J_{R}(\pi)$
$s.t.\quad J_{C}(\pi)\leq\epsilon$
where $J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, $J_{C}(\pi)=c(s_{t},a_{t}),\forall t\in[0,\infty)$, $\pi(s_{t},a_{t})$ is the live stream allocation policy to be optimized.
A common method to solve the above CMDP is to transform the problem into a min-max optimization by introducing the Lagrange multiplier $\lambda$:
(3) $(\pi^{*},\lambda^{*})=\arg\min_{\lambda\geq 0}\max_{\pi}J_{R}(\pi)-\lambda(J_{C}(\pi)-\epsilon)$
Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function by substituting into $r(s_{t},a_{t})$ and $c(s_{t},a_{t})$ of Eq. 1:
(4) $\hat{r}(\lambda,s_{t},a_{t})$ $=r(s_{t},a_{t})-\lambda c(s_{t},a_{t})$
$=y_{l}-\lambda(\frac{1}{B}y_{v}-y_{l})$
$=(1+\lambda)y_{l}-\frac{\lambda}{B}y_{v}$
In practice, we solve the unconstrained MDP problem with a simplified reward function as follows:
(5) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(y_{l}-\frac{\lambda}{B}y_{v}\right)\right]$
where $\lambda$ is a hyper-parameter to carefully control the final live stream allocation ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint. Note that we leave the further discussion of the rationality of the above reward function in the Appendix C.
3. Proposed Framework
To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig. 2. SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules. For clarity, only one critic network is shown in Fig. 2, while the remaining three critic networks, which use the same architecture, are omitted.
3.1. User & live stream Feature Extraction Module
The user & live stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$. First, we use a unified embedding layer to map the user’s static features, live stream features, historical live stream list features, and short video-list features into low-dimensional dense representations. We denote $\mathbf{v}_{U}$, $\mathbf{v}_{L}$, $\left\{\mathbf{e}_{1},\cdots,\mathbf{e}_{M_{l}}\right\}$, $\left\{\mathbf{e}^{\prime}_{1},\cdots,\mathbf{e}^{\prime}_{M_{v}}\right\}$ as the corresponding embedding vectors or sets of embedding vectors, respectively. We then define $\mathbf{h}_{a}=[\mathbf{v}_{U},\mathbf{v}_{L}]$ as the concatenation of $\mathbf{v}_{U}$ and $\mathbf{v}_{L}$.
To aggregate historical live stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
(6) $\mathbf{h}_{live}$ $=\sum_{i=0}^{M_{l}}f_{l}(\mathbf{h}_{a},\mathbf{e}_{i})\mathbf{e}_{i}$
$\mathbf{h}_{video}$ $=\sum_{j=0}^{M_{v}}f_{v}(\mathbf{h}_{a},\mathbf{e}^{\prime}_{j})\mathbf{e}_{j}^{\prime}$
where $f_{l}$, $f_{v}$ are different target attention functions, such as a feed-forward network whose output is a normalized score.
After aggregating the historical live stream and short video features via two separate attention networks, we concatenate all of the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$. We then use a shared multi-layer perceptron (MLP) for both the actor and critic networks to obtain the latent representation of state $s_{t}$, i.e. $\mathbf{h}^{\prime}_{s}=f_{MLP}(\mathbf{h}_{s})$. To ensure training stability, we stop the gradient flow through $\mathbf{h}^{\prime}_{s}$ for the actor network and only use the more complex critic networks to optimize $\mathbf{h}^{\prime}_{s}$, as we find that sharing the feature extraction module and the $f_{MLP}$ network causes interference between the policy and critic gradients.
3.2. Multi-Group State Decomposition Module
Since live stream content is not equally appealing to all users of the short video app Kwai, it is crucial to inject live streams selectively into the short video feed. Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream allocation policy for each individual user. Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
We find that prior information on diverse user groups is essential to improve the accuracy of the decision making of RL models in our scenario, which provides additional user preference information to the policy. A natural approach to distinguish between different user behaviors is to partition users into distinct groups based on their historical activity. Specifically, we categorize users into $K$ disjoint groups according to their activity level of the live stream, which is determined by the cumulative watch time of the live stream during the past 3 weeks. Users with a higher historical watch time are assigned to groups with higher activity levels.
Empirically, we show the effectiveness of this module in enhancing the stability of the RL model in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
3.3. Supervised Learning-enhanced Actor Critic
To alleviate the impact of drastic changes in data distribution that could lead to the divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from getting trapped in model collapse due to large cumulative errors in the critic learning process.
3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
From the proof in Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant, and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018). This theoretically indicates that a network with Layer Normalization is less sensitive to input variations and can maintain stable gradients despite perturbations during model training. Hence, we apply Layer Normalization to the inputs of both the actor and critic networks, as shown in Fig. 2.
3.3.2. Critic Network
Let $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})\in\mathcal{D}$ be a training sample from our real-time dataset $\mathcal{D}$. To address the maximization bias problem (Van Hasselt et al., 2016), we employ four critic networks: two current Q-networks $Q_{\phi_{1}}$, $Q_{\phi_{2}}$ and two corresponding target Q-networks $Q^{\prime}_{\phi_{1}}$, $Q^{\prime}_{\phi_{2}}$ in Clipped Double Q-Learning (Fujimoto et al., 2018). The critic learning objective is as follows:
(7) $\mathcal{L}_{Critic}=\sum_{i=1}^{2}\mathbb{E}_{(s,a)\in\mathcal{D}}$ $\left[Q_{\phi_{i}}(s_{t},a_{t})-Q_{label}(s_{t+1})\right]^{2}$
$Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\max_{a_{t+1}}\hat{Q}^{\prime}(s_{t+1},a_{t+1})$
where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use Huber loss (Hastie et al., 2009) to optimize the above objective.
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks. Specifically, due to the instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term. Then the $r(s_{t},a_{t})$ term could not provide any information to guide the critic learning. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we divide the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
(8) $Q_{\phi_{i}}(s_{t},a_{t})$ $=R_{\theta_{i}}(s_{t},a_{t})+\gamma T_{\xi_{i}}(s_{t},a_{t}),\quad i=1,2$
$Q^{\prime}_{\phi_{i}}(s_{t+1},a_{t+1})$ $=R^{\prime}_{\theta_{i}}(s_{t+1},a_{t+1})+\gamma T^{\prime}_{\xi_{i}}(s_{t+1},a_{t+1}),\quad i=1,2$
where $R_{\theta_{i}}$ and $T_{\xi_{i}}$ are RPN and QRN, $R^{\prime}_{\theta_{i}}$ and $T^{\prime}_{\xi_{i}}$ are target RPN and QRN, respectively. We use separate MLPs to model the RPN and QRN.
Since the distribution of live stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq. 5 becomes difficult to learn. In this work, we propose a distribution discretization method to improve reward learning.
Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time. Let $y$ denote the actual live stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
(9) $y$ $=y_{st}+\delta\cdot(y_{end}-y_{st})$
which resembles the structure of linear reward shifting (Sun et al., 2022).
Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls. Note that a separate bin is set for the case $a_{t}=0$. Then, the RPN $R_{\theta_{i}}$ can be modeled by multi-task neural networks:
(10) $R_{\theta_{i}}(s_{t},a_{t})$ $=sigmoid\left(F_{\Gamma_{i}}(s_{t},a_{t})-G_{\Theta_{i}}(s_{t},a_{t})\right)$
$F_{\Gamma_{i}}(s_{t},a_{t})$ $={o(t)_{l}}^{T}\left(y_{l}^{st}+f_{\Gamma_{i}}(s_{t},a_{t})\cdot\left(y_{l}^{end}-y_{l}^{st}\right)\right)$
$G_{\Theta_{i}}(s_{t},a_{t})$ $={o(t)_{v}}^{T}\left(y_{v}^{st}+g_{\Theta_{i}}(s_{t},a_{t})\cdot\left(y_{v}^{end}-y_{v}^{st}\right)\right)$
where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin. Note that we introduce the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq. 10 will not affect the calculation of $Q^{\prime}_{\phi_{i}}(s_{t+1},\cdot)$, because we have already reserved $r_{t+1}$ in the dataset $\mathcal{D}$. Therefore, $o(t+1)_{l}$ and $o(t+1)_{v}$ can be easily obtained from $r_{t+1}$ to compute $R^{\prime}_{\theta_{i}}(s_{t+1},\cdot)$.
As shown in Eq. 10, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin. We then reconstruct the predicted live streaming and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively. Finally, we obtain the predicted normalized reward, $R_{\theta_{i}}(s_{t},a_{t})$. To optimize the predicted watch-time proportions, we use Huber loss as follows:
(11) $\mathcal{L}_{SL}=\sum_{i=1}^{2}Huber\_Loss(\delta_{l},F^{\prime}_{\Gamma_{i}})+Huber\_Loss(\delta_{v},G^{\prime}_{\Theta_{i}})$
where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live stream and short video watch times, respectively.
The reason for using the Huber loss on the time ratio $\delta$ instead of the actual watch time $y$ is that the time ratio is within the range $[0,1]$, resulting in smaller gradients in the neural network. Furthermore, the variance of $\delta$ is much smaller than that of the original watch time $y$. Consequently, the outputs of the learned reward networks ($f_{\Gamma_{i}}$ and $g_{\Theta_{i}}$) will exhibit smaller variances. Hence, reward learning with the discretization of watch time distributions can be viewed as a novel variance reduction technique.
3.3.3. Actor Network
We also incorporate the multi-group state decomposition module into the actor network. The loss function of the actor network is shown below:
(12) $\mathcal{L}_{Actor}$ $=\mathbb{E}_{(s,a)\in\mathcal{D}}\left[-\hat{Q}(s_{t},a_{t})\log p(s_{t},a_{t})\right]$
where $\hat{Q}(s_{t},a_{t})=\min_{i=1,2}Q_{\phi_{i}}(s_{t},a_{t})$, and $p(s_{t},a_{t})$ is the action probability output of the actor network.
Moreover, we observe that high values of $\hat{Q}(s_{t},a_{t})$ can cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$. To address this, we apply softmax normalization to $\hat{Q}(s_{t},a_{t})$ to obtain $\hat{Q}_{norm}=softmax(\hat{Q}(s_{t},a_{t}))$, and propose a modified actor loss function:
(13) $\mathcal{L}_{Actor}$ $=\mathbb{E}_{(s,a)\in\mathcal{D}}\left[-\hat{Q}_{norm}(s_{t},a_{t})\log p(s_{t},a_{t})\right]$
A similar loss function can be found in AWAC (Nair et al., 2020), and its theoretical results demonstrate that the above loss is equivalent to an implicitly constrained RL problem:
(14) $\pi_{k+1}$ $=\arg\max_{\pi\in\Pi}\mathbb{E}_{\mathbf{a}\sim\pi(\cdot|\mathbf{s})}\left[\hat{Q}^{\pi_{k}}(s_{t},a_{t})\right]$
${\rm s.t.}$ $D_{\rm KL}\left(\pi(\cdot|\mathbf{s})\|\pi_{o}(\cdot|\mathbf{s})\right)\leq\epsilon$
where $D_{\rm KL}$ is the Kullback-Leibler (KL) divergence, and $\pi_{o}$ is the behavior policy derived from the dataset $\mathcal{D}$.
We note that the actor loss in Eq. LABEL:eq:15 is a standard cross-entropy loss function. From the perspective of knowledge distillation (Hinton, 2015), the actor distills policy knowledge from more complex critic networks. The teacher critic networks guide the more lightweight student actor network to adjust and converge to an optimal policy. In practice, we only need to deploy the actor network in the online live stream recommendation system, which significantly reduces computational complexity and improves real-time response speed.
Overall, the final loss function of our proposed SL-MGAC algorithm is as follows:
(15) $\mathcal{L}=\mathcal{L}_{Actor}+\mathcal{L}_{Critic}+\mathcal{L}_{SL}$
3.4. Online Exploration and Deployment
For online exploration, we adopt the commonly used $\epsilon$-greedy (Watkins, 1989) strategy:
(16) $\pi_{online}(s_{t})=\begin{cases}{\rm random\;action\;from\;}\mathcal{A}(s_{t}),\quad&if\;\psi<\epsilon\\ \arg\max_{a\in\mathcal{A}(s_{t})}\pi(a_{t}|s_{t}),&{\rm otherwise}\end{cases}$
where $\psi$ is a random number, and $\epsilon$ is maximal exploration probability.
Figure 3. System Architecture of the SL-MGAC algorithm.
We implemented the SL-MGAC algorithm in our recommendation system, and the overall architecture of the system is shown in Fig. 3. The online RL agent collects real-time user interaction logs, while the offline model trainer optimizes the SL-MGAC model in an off-policy manner using streaming training data. Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
4. Experiments
We perform both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live streaming allocations in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio. In addition, the online environment contains more than 100 million app users for AB-Tests.
4.2. Compared Methods
4.2.1. Baselines
We compare our approach with existing non-reinforcement learning and reinforcement learning methods.
  • Learning to Rank (L2R) (Chen and Zhou, 2020). A supervised learning framework that predicts the reward for each action. The action with the maximum reward is selected as the agent’s action.
  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of updating the target network.
  • BCQ (Fujimoto et al., 2019). A widely used offline reinforcement learning algorithm that adds action restrictions and encourages the agent to behave similarly to the behavior policy.
  • SAC (Haarnoja et al., 2018; Christodoulou, [n. d.]). A classic off-policy reinforcement learning method that maximizes the trade-off between cumulative reward and policy entropy. We use the discrete version of SAC in later experiments.
  • TD3 (Fujimoto et al., 2018). A modified version of DDPG (Lillicrap, 2015) addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
  • TD3-BC (Fujimoto and Gu, 2021). An offline reinforcement learning variant of the TD3 algorithm, with behavior cloning (BC) regularization to constrain policy learning.
  • IQL (Kostrikov et al., 2021). An offline reinforcement learning algorithm that leverages expectile regression in Q-Learning.
Note that in our optimal live stream allocation control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
4.2.2. Variations of our model
We also compare the SL-MGAC algorithm with several variants in an ablation study to illustrate the effectiveness of multi-group state decomposition, supervised learning for critic learning, distribution discretization (DD), and other techniques.
  • SL-MGAC (w/o MG): An SL-MGAC variant without the Multi-Group State Decomposition module.
  • SL-MGAC (w/o MG & DD): An SL-MGAC variant without the Multi-Group State Decomposition module and the distribution discretization technique in reward learning.
  • SL-MGAC (w/o MG & DD & SL): An SL-MGAC variant without the Multi-Group State Decomposition module, the distribution discretization technique in reward learning, and the supervised learning procedure.
  • SL-MGAC (w/o LN): An SL-MGAC variant without the Layer Normalization technique.
  • SL-MGAC (w/o SG): An SL-MGAC variant without the Stop Gradient technique for hidden state representation in the actor network.
  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q-value normalization in actor loss.
  • SL-MGAC-0: An SL-MGAC variant with the discount factor $\gamma=0$.
  • SL-MGAC-sep: An SL-MGAC variant with separate optimization of the actor network. Hence, $\hat{Q}_{norm}$ in Eq. LABEL:eq:15 is detached from the computation graph.
  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the loss of the critic, as shown below, where $p(s_{t+1},\cdot)$ is the probability of action of the output of $s_{t+1}$ by the actor network.
    (17) $Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\sum_{a_{t+1}}p(s_{t+1},a_{t+1})\hat{Q}(s_{t+1},a_{t+1})$
4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
4.4. Offline Policy Evaluation
We follow the approach in (Xue et al., 2023) and adopt a commonly used offline policy evaluation method, namely Normalized Capped Importance Sampling (NCIS) (Swaminathan and Joachims, 2015), to evaluate performance. The evaluation metric is the cumulative reward across all trajectories of test users.
Methods Cumulative Reward
L2R (Chen and Zhou, 2020) 413.49
DQN (Mnih et al., 2015) 413.18
BCQ (Fujimoto et al., 2019) 416.69
SAC (Christodoulou, [n. d.]) 435.41
TD3 (Fujimoto et al., 2018) 432.08
TD3-BC (Fujimoto and Gu, 2021) 430.23
IQL (Kostrikov et al., 2021) 432.03
SL-MGAC-sep 435.19
SL-MGAC-vanilla 450.21
SL-MGAC 458.49
Table 1. Overall offline performance of compared methods.
The results of the offline policy evaluation are shown in Table 1. Compared to the supervised learning method L2R, most RL methods achieve higher cumulative rewards, except for DQN. SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live stream & short video mixed recommendation system with drastically changing data distributions. Furthermore, SL-MGAC outperforms SL-MGAC-vanilla, suggesting that incorporating the actor network in the Q label calculation of Eq. 17 may interfere with the learning of the critic to some extent, slightly affecting the performance of the SL-MGAC-vanilla.
4.5. Ablation Study
We compare the offline performance of different SL-MGAC variants. As shown in Table 2, SL-MGAC outperforms all other variants on our offline dataset, demonstrating the effectiveness of the proposed multi-group state decomposition module, supervised learning procedure, and other techniques. Moreover, we find that SL-MGAC (w/o Q-norm) achieves the lowest reward, indicating that the Q-normalization technique in the actor loss enhances the model’s convergence and improves its performance.
Methods Cumulative Reward
SL-MGAC (w/o MG) 454.62
SL-MGAC (w/o MG & DD) 452.59
SL-MGAC (w/o MG & DD & SL) 449.13
SL-MGAC (w/o LN) 453.61
SL-MGAC (w/o SG) 457.26
SL-MGAC (w/o Q-norm) 392.12
SL-MGAC 458.49
Table 2. Offline performance of SL-MGAC variants.
Next, we compare the training processes of SL-MGAC and SL-MGAC (w/o MG), as shown in Fig. 4. We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller. This demonstrates the effectiveness and variance reduction effect of the MG-SD module.
Figure 4. The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training. The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations.
4.6. Parameter Sensitivity
We analyze the impact of the number of user groups $K$ on the performance of SL-MGAC. As shown in Fig. 5, the proposed SL-MGAC algorithm with $K=6$ achieves the highest cumulative reward, demonstrating that increasing the number of user groups within a certain range improves model performance.
Figure 5. Performance of different numbers of user group $K$.
Methods live stream DAU live stream Watch Time Video Watch Time App Usage Duration User Retention
L2R +0.321% -0.558% -0.238% -0.178% -0.154%
Dummy +51.3% +34.6% -2.216% -1.062% -0.439%
SL-MGAC-0 +3.928% -2.934% -0.153% -0.117% -0.102%
SL-MGAC +2.616% +7.431% +0.197% +0.121% +0.086%
Table 3. Online A/B Test Performance of compared methods.
4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline. The baseline uses the SAC framework for live stream allocation. The results of the online A / B test are shown in Table 3.
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves a significant improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. This shows that SL-MGAC is more effective in maximizing long-term live stream rewards while avoiding the negative impact on app usage duration and user retention. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users. Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage time and user retention.
4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 6, the live stream allocation ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream allocation ratio refers to the proportion of requests with action $a=1$.
Figure 6. The trend curve of online live stream allocation ratio during a whole day.
We calculate the amplitudes of the live stream allocation ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 8 in Appendix B. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
5. Related Work
5.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize cumulative long-term rewards over time, which has attracted significant attention in the research of recommendation systems in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend complete item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions. CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving the user experience. Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
A similar approach, called self-supervised actor-critic (Xin et al., 2020), combines supervised learning with critic learning. Their supervised learning task focuses on predicting the next item recommendation probability, whereas our approach introduces supervised learning to restrict critic learning.
5.2. RL divergence and instability
Recently, there has been a growing literature that focuses on RL divergence and instability. (Van Hasselt et al., 2016) introduces Double DQN to mitigate the maximization bias problem in DQN. (Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q values. (Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve the stability of training. (Dohare et al., 2023) investigates policy collapse in a 2-state MDP and finds that L2 regularization and the non-stationary Adam optimizer (Kingma, 2014) are both effective in alleviating RL instability. (Yue et al., 2024) theoretically analyzes the causes of RL divergence and applies Layer Normalization to mitigate RL divergence and instability.
6. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention in the challenging context of a live stream & short video mixed recommendation system. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model. In practice, we minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully implemented in large-scale industrial recommendation systems. Currently, we are trying to apply the SL-MGAC methods to other similar recommendation scenarios, including E-commerce and advertising. We also aim to continuously improve the performance of SL-MGAC in further research. Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
7. Acknowledgment
The authors acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
Appendix A Tables
Hyper-parameter Value
MGN in Actor [128, 63, 31, 2]
MGN in RPN [128, 64, 32, 8]
MGN in QRN [128, 64, 32, 2]
Live-Time DDB [0s, 6s, 15s, 30s, 60s, 100s, 600s, 1200s]
Video-Time DDB [0s, 3s, 10s, 25s, 50s, 100s, 600s, 1200s]
Optimizer Adam
User Group Number $K$ 6
Online Exploration $\epsilon$ 0.2
Table 4. Hyper-parameters of SL-MGAC.
We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, and DDB as Distribution Discretization Bins. Both live and video watch time distribution have 7 time bins. We leave a separate time bin for the $a=1$ case, and hence the output layer dimension of RPN is 8.
Appendix B Figures
Figure 7. Data distributions of online live stream room (or viewer) count, live stream watch time and short video watch time.
Figure 8. The amplitude distribution of live stream injection ratio.
Appendix C More Ablation Studies